# Exploring Missing Modality in Multimodal Egocentric Datasets

### v1

@@ 4 Experiments
We present a detailed analysis of our MMT under both training-with-complete and training-with-missing. For both scenarios, we explore both strategies fixed-replace and random-replace to train MMT (Sections 4.34.4). We report classification accuracy under several missing rates $r_{test}$ for three egocentric datasets. Additionally, we study the effect of fusion layers $L_{f}$ in Section 4.5. Then, we compare our method vs. a baseline proposed by [25] in Section 4.6.
@@ 4.1 Datasets
Epic-Sounds [18] spans the same 100 video hours as Epic-Kitchens but is annotated with sound labels. This dataset does not follow the noun and verb annotations from Epic-Kitchens, instead, it has 44 unique class labels. The dataset is composed of 79K annotated clips.
@@
Table 2: The performance of the audio, video, and bottleneck audiovisual models on each dataset. We train and evaluate all models with $r_{train}$ = 0. In all datasets, multimodal performance beats unimodal performance.
Figure 3: Missing modality rate in training data $r_{train}$ vs. accuracy when training with fixed-replace strategy. We show that our MMT is effective in learning with missing modalities under several $r_{train}$ regimes. For Epic-Sounds and Ego4d-AR the higher masking seems to be more beneficial to learn MMT. However, for Epic-Kitchens lower masking yields better accuracy overall.
In Table 2, we show the unimodal and multimodal performance for each downstream dataset. As expected from the annotation strategy, video is the dominant modality in Epic-Kitchens and Ego4d-AR, and audio is the one in Epic-Sounds. Therefore, as mentioned in Section 3.1, we train our models to be robust to missing video in Epic-Kitchens and Ego4D-AR and to missing audio in Epic-Sounds. Since Epic-Sounds has fewer training samples and has a more balanced unimodal performance of each modality, we use it more extensively in our analysis.
For training-with-complete, we use Epic-Sounds and Epic-Kitchens. For training-with-missing, we use Ego4D-AR and create a modal-incomplete version of Epic-Sounds by enforcing $r_{train}>0$.
4.2 Implementations Details
Architecture. Following [17], we use ViT-Base [6] with 12 transformer layers, 12 attention heads, and embedding dimension 768 as the encoder for each modality. For the fusion design, we follow MBT [32] and fix the number of bottlenecks to $B=4$ and the fusion layer to $L_{f}=8$ (except for the fusion layer ablation).
Making training sets with different fixed $r_{train}$. We begin by taking all training instances that are complete across all modalities and shuffle them randomly once. For the Ego4D-AR dataset, instances that are inherently missing are placed at the very start of this sequence. Following this, we establish our desired missing rate, represented by $r_{train}$ by sampling from the start of this list. This method ensures that any increase in $r_{train}$ builds upon the existing set of modal-incomplete instances, meaning that if we increase $r_{train}$, we add new modal-incomplete instances to those already included, thereby maintaining a cumulative effect.
4.3 The effect of the number of modal-incomplete samples $r_{train}$ in fixed-replace.
We analyze the effect of $r_{train}$, the proportion of the modal-incomplete inputs in the training data, under the fixed-replace strategy in each dataset. For Epic-Kitchens we use $r_{train}\in\{25\%,50\%,75\%\}$, and $r_{train}\in\{29\%,50\%,75\%\}$ for Ego4D-AR. Recall that Ego4D-AR has a natural $r_{train}=29\%$, and we can’t reduce this value. We observed that higher values of $r_{train}$ yield better performance in Epic-Sounds; hence, we use $r_{train}\in\{40\%,60\%,80\%\}$ in this dataset.
Figure 3 shows the results for each dataset. We can see how Epic-Sounds and Ego4D-AR follow a similar trend: with the lowest $r_{train}$ in each dataset, the model does not adapt well to the missing inputs. For example, at $r_{test}=100\%$ in Ego4D-AR, the model trained with $r_{train}=50\%$ is $8\%$ points more accurate than the model trained with $r_{train}=29\%$. Furthermore, training with higher $r_{train}=50\%$ and $r_{train}=75\%$ still yields reasonable accuracy at lower $r_{test}$. The opposite is true for Epic-Kitchens, where not only does the model trained with the lowest $r_{train}=25\%$ adapt well to the missing modality, but the accuracy drops significantly when we train with more modal-incomplete inputs. Overall, we highlight that: (1) for all datasets, higher $r_{train}$ results in higher accuracy on the fully modal-incomplete test set ($r_{test}=100\%$), emphasizing the importance of training MMT with more samples; (2) surprisingly, while higher $r_{train}$ causes more modality information loss, we can still train accurate models in Epic-Sounds and Ego4D-AR.
Figure 4: Modality drop probability $p$ vs. accuracy for modal-complete Epic-Sounds and Epic-Kitchens. In Epic-Sounds, as $p$ decreases to $40\%$ the model’s performance with missing audio goes down. While dropping with high $p=80\%$, the model still reaches decent performance at all $r_{test}$. In Epic-Kitchens, we observe very similar behavior as in fixed-replace.
4.4 The effect of the Modality Drop Probability $p$ in random-replace
We analyze the impact of $p$, the probability with which the tokens of modal-complete inputs are replaced with MMT. For easier comparison with the results in 4.3, we use $p\in\{25\%,50\%,75\%\}$ for Epic-Kitchens and Ego4D-AR, and $p\in\{40\%,60\%,80\%\}$ for Epic-Sounds.
Training-with-complete Figure 4 shows changes in performance with different $p$ in training-with-complete. We observe how $p$ exhibits a similar effect in random-replace as $r_{train}$ in fixed-replace. However, there are some noticeable differences. First, across all $p$ values, the models trained with random-replace reach the same or higher accuracy compared to the corresponding models trained with fixed-replace (Figure 3). This is especially evident in Epic-Sounds: by switching from fixed-replace to random-replace, we get an average improvement of $2.3\%$, $1.6\%$, and $1.4\%$ points for $p=r_{train}$ of $40\%$, $60\%$, and $80\%$, respectively. We attribute the greater improvement with the lower $p=40\%$ to the ability of the model to ”see” more of the same samples appear in both missing and complete modality scenarios, as we outline in 3.4. When $p$ is too high, fewer samples have a chance of being seen by the model as modal-complete multiple times during training. Furthermore, while the improvement is the lowest for $p=80\%$, this model still performs optimally at $r_{test}=0\%$, which is not the case for the model with $r_{train}=80\%$. Thus, we find random-replace more effective for training MMT.
Training-with-missing In Figure 5, we show the results of training with random-replace in training-with-missing scenario. Note how we are using both $r_{train}$ and $p$ here. Because of that, the numbers are hard to compare to other results directly, but we can still see similar patterns. Consistently with all previous results, the more samples we replace with MMT during training, the better the model adapts to severely missing modality at test time. Furthermore, dropping the modality of too many inputs (randomly or fixed) negatively affects the performance. For example, while the model trained with $p=75\%$ achieves higher accuracy at $r_{test}=100\%$ in Ego4D-AR, it seems that it learns to ignore the audio completely, as the performance at $r_{test}=0\%$ drops to the unimodal video performance in this dataset. Additionally, by distilling the results in Figures 3 and 5, we observe that using random-replace in Epic-Sounds with $r_{train}>0$ is more effective for training with missing modalities.
@@ 4.4 The effect of the Modality Drop Probability $p$ in random-replace
Figure 5: Results with the modal-incomplete training data. As Epic-Sounds does not naturally have missing modality in the training data, we manually remove the audio from $r_{train}=25\%$ = and $r_{train}=50\%$ of samples in the train set. In Ego4D-AR, there is a naturally missing audio in $r_{train}=29\%$ of the samples.
@@
Figure 6: Fusion Layer $L_{f}$ vs. accuracy in the models trained with no adaptation strategy (Non-Adapted) and trained with MMT (Adapted). For training with MMT, we use random-drop strategy with $p=60\%$ for Epic-Sounds and $p=25\%$ for Epic-Kitchens. This strategy makes MBT more robust to missing modalities across all $L_{f}$ and significantly reduces the negative effect of missing modalities.
@@
As we mentioned in Sec. 3.2, the fusion layer does affect the performance of the bottleneck model, especially when test inputs are modal-incomplete. In Fig. 6, we examine whether training with Modality Masking Training (MMT) enhances robustness to a missing modality at various fusion layers and whether models trained with MMT (adapted) show the same sensitivity to the fusion layer as those trained without it (non-adapted). As we observe, the introduction of MMTs makes the model more robust to the missing modality across all fusion layers in both Epic-Sounds and Epic-Kitchens. With extremely severe $r_{test}=100\%$, the adapted models perform with $\sim 45\%$ accuracy in Epic-Kitchens, while the unimodal audio model achieves $40\%$ in this dataset. Furthermore, these adapted models do not exhibit similar sensitivity to the fusion layer as non-adapted. For example, in Epic-Sounds, the non-adapted models trained with $L_{f}=11$ exhibit superior performance, which is not the case for the adapted models (at $r_{test}=100\%$, $37\%$ accuracy with $L_{f}=11$ but $\sim 40\%$ for other $L_{f}$). Overall, all adapted models perform consistently well, each providing decent performance in a modal-incomplete inference. Our approach effectively addresses a long-standing issue of selecting the appropriate fusion layer when faced with missing modalities [31].
4.6 MMT vs. other methods
@@
In Tables 6(a)6(b) and 6(c), we report the test results using our fixed-replace and random-replace strategies and the baseline method for Epic-Sounds, Epic-Kitchens, and Ego4D-AR, respectively. We also report the accuracy of the unimodal model in each dataset. For the baseline, we implement the multimodal prompts method [25]. It was originally based on ViLT [23] for image-text classification, so we implemented our version for action recognition based on MBT. Note that this method relies on a strong pre-trained backbone to efficiently finetune it by optimizing a few network parameters. As we deal with audiovisual learning in egocentric videos, we rely on large-scale egocentric pre-training using MAEs. As seen in Tables 6(a)6(b)6(c), our strategy outperform the alternative proposed by [25].
$r_{test}$ 0% 25% 50% 75% 100%
unimodal 41.4 41.4 41.4 41.4 41.4
no adaptation 55.2 45.6 37.1 28.3 19.5
mm prompts [25] 36.7 33.3 30.0 26.1 22.6
fixed-replace(ours) 55.3 51.2 47.6 43.4 40.2
random-replace(ours) 56.3 52.3 48.6 44.6 40.7
$r_{test}$ 0% 25% 50% 75% 100%
unimodal 40.0 40.0 40.0 40.0 40.0
no adaptation 63.9 55.5 46.8 37.9 29.5
mm prompts [25] 32.5 31.1 30.2 29.2 28.2
fixed-replace (ours) 63.4 58.8 54.1 49.5 44.8
random-replace (ours) 63.4 59.0 54.6 50.0 45.3
$r_{test}$ 27% 50% 75% 100%
unimodal 34.6 34.6 34.6 34.6
no adaptation 30.5 25.9 21.1 16.0
mm prompts [25] 18.4 16.4 13.9 11.5
fixed-replace (ours) 36.2 34.9 33.6 32.4
random-replace (ours) 36.7 35.2 33.5 32.0
@@
Table 3: Comparison of our method with baselines on various datasets. We demonstrate the accuracy of our method vs. the adaptation of the method proposed by [25] and unimodal performance across different missing modality ratios $r_{test}$. We show in bold the best result and underline the runner-up.
@@
We find that across all datasets, our fixed-replace and random-replace strategies either reach (in Epic-Sounds and Ego4D-AR) or exceed (in Epic-Kitchens) the unimodal performance in extreme $r_{test}=100\%$. Interestingly, in Epic-Sounds, random-replace also regularizes the training and increases the $r_{test}=0\%$ performance (Table 6(a)) by 1.1%.
@@ 5 Conclusion
We explore the missing modality problem in multimodal egocentric datasets. We suggest a simple yet effective method by learning the optimal token representation of the missing modality (MMT). Placing learnable tokens to represent missing inputs provides an easy and intuitive way to train and test with modal-incomplete inputs. We propose 2 ways to learn MMT when training action recognition models and show how their performance brings us closer to robust and effective multimodal systems.

### v2

@@ 4 Experiments
We present a detailed analysis of our MMT under both modal-complete and modal-incomplete training sets. For both scenarios, we explore the usage of random-replace. Namely, in Sec. 4.4 we ablate the effect of $p$. In Sec. 4.5, we study how the severity of missing modality in the training data affects the performance. Sec. 4.6 we extend the setup to multiple missing modalities. Additionally, we study the effect of fusion layers $L_{f}$ in Sec. 4.7. Then, we compare our method to the baselines we proposed and [30] in Sec. 4.8.
@@ 4.1 Datasets
Epic-Sounds [22] spans the same 100 video hours as Epic-Kitchens but is annotated with sound labels. This dataset does not follow the noun and verb annotations from Epic-Kitchens; instead, it has 44 unique class labels. The dataset is composed of 79K annotated clips.
@@
4.2 Implementations details
Architecture. Following [21], we use ViT-Base [6] with 12 transformer layers, 12 attention heads, and embedding dimension 768 as the encoder for each modality. For the fusion design, we follow MBT [37] and fix the number of bottlenecks to $B=4$ and the fusion layer to $L_{f}=8$ (except for Sec. 4.7).
Table 1: The performance of the audio, video, and bottleneck audiovisual models on each dataset. We train and evaluate all models with $r_{train}$ = 0 and $r_{test}=0\%$. In Ego4D-AR, this is done by filtering out the modal-incomplete samples. In all datasets, multimodal performance beats unimodal performance.
@@ 4.2 Implementations details
4.3 Unimodal and baseline multimodal models
In Table 1, we show the unimodal and multimodal performance for each downstream dataset. As the original MBT, the multimodal models are trained with fully modal-complete samples. For our baseline results across all datasets, we employ these multimodal models. In scenarios where $r_{test}$ is small, indicating that the testing data is nearly modal-complete, it is desired that the adapted model exhibits performance more closely aligned with the multimodal baseline model. Consequently, this adaptation should uphold the multimodal reasoning capabilities of the model.
As expected due to the annotation strategy, video serves as the dominant modality in Epic-Kitchens and Ego4d-AR, and audio takes precedence in Epic-Sounds. Consequently, as detailed in Sec. 3.1, we train our models to be robust to missing video in Epic-Kitchens and Ego4D-AR and to missing audio in Epic-Sounds (except for Sec. 4.6 where we assume that any modality could be absent). Additionally, we report the unimodal performance of the non-missing modality (referred to as unimodal, i.e., video in Epic-Sounds) as we expect the adapted models to converge towards this performance at higher values of $r_{test}$.
Since Epic-Sounds has fewer training samples and exhibits a more balanced unimodal performance across each modality, we leverage it more extensively in Sec. 4.5 and Sec. 4.6.
4.4 Results with MMT.
Figure 3: Modality drop probability $p$ vs. accuracy for modal-complete Epic-Sounds, Epic-Kitchens and Ego4D-AR. In all datasets, our method dramatically improves the performance of the baseline (orange).
We apply MMT as mentioned in Sec. 3.4. Fig. 3 shows changes in performance with different $p$ in random-replace and compares it with the unimodal (purple) and baseline multimodal (orange) performance in each dataset. We use $p\in\{12.5\%,25\%,50\%\}$ for Epic-Kitchens. Ego4D-AR has naturally missing modality with $r_{train}=29\%$, and we ablate with $p\in\{0\%,25\%,50\%,75\%\}$. We observed that higher values of $p$ yield better performance in Epic-Sounds; hence, we use $p\in\{30\%,60\%,90\%\}$ in this dataset.
By looking at the model’s performance across all $p$ values, we notice the importance of picking $p$ large enough for the model to adapt well but not too large to avoid high information loss. On the one hand, if there are insufficient training samples for MMT (smaller $p$), the model does not perform optimally at higher $r_{test}$. For example, in Epic-Sounds, the model trained with $p=30\%$ reaches 42.7% at $r_{test}=25\%$ while increasing $p$ to 60% gives 44.6%. Similarly, if we only use naturally modal-incomplete samples to train MMT in Ego4D-AR ($p=0\%$), the model reaches 28.6% at $r_{test}=25\%$, while if we increase to $p=25\%$, the performance increases to 33.5%. On the other hand, if $p$ is too large, the model performs suboptimally for more modal-complete test data (smaller $r_{test}$). For example, while the model trained with $p=75\%$ achieves higher accuracy at $r_{test}=100\%$ in Ego4D-AR, it seems that it learns to ignore the audio completely, as the performance at $r_{test}=0\%$ drops to the unimodal video performance in this dataset.
We find that the models trained with $p=60\%$, $p=25\%$, and $p=25\%$ in Epic-Sounds, Epic-Kitchen, and Ego4D-AR, respectively, yield the best overall performance in each dataset. These models perform significantly better than the baseline models (orange) in the missing modality scenarios. For instance, at $r_{test}=50\%$, the adapted models improve the baseline performance by $11.5\%$ points in Epic-Sounds, $7.8\%$ points in Epic-Kitchens, and $9.3\%$ points in Ego4D-AR. Furthermore, for Epic-Sound and Epic-Kitchens, the models trained with MMT reach or surpass the unimodal performance at extremely severe $r_{test}=100\%$. Nevertheless, the adapted models maintain the multimodal baseline performance at $r_{test}=0\%$, showing that the adaptation strategy does not harm the model’s capabilities.
Note how in Ego4D-AR, the baseline fails to reach the unimodal accuracy event at the lowest $r_{test}=27\%$, making our models trained with MMT a significantly better choice. This happens because $r_{train}=29\%$ of samples were filtered out to train the multimodal baseline, causing information loss. This demonstrates how MMT enables better leverage of data in modal-incomplete datasets.
In Epic-Kitchens, the model performs better with $p=25\%$ much smaller than in Epic-Sounds. We believe this is because video (the missing modality) produces almost 4×4$\times$4 × more tokens than audio; thus, replacing video tokens causes more information loss.
@@ 4.3 Unimodal and baseline multimodal models
Figure 4: Results with the modal-incomplete training data. As Epic-Sounds does not naturally have missing modality in the training data, we manually remove the audio from (left) $r_{train}=25\%$ and (right) $r_{train}=50\%$ of samples in the train set.
4.5 Datasets with modal-incomplete training data.
In Sec. 4.4, Ego4D-AR is the only naturally modal-incomplete dataset, and we want to know how our method generalizes in other datasets with different $r_{train}.$ What happens if a dataset has even higher $r_{train}$ than Ego4D-AR? Filtering out noisy or corrupted data causes significant information loss if $r_{train}$ is high. By using MMT, we can mitigate that and learn from all instances.
We show the results with our modal-incomplete version of Epic-Sounds with $r_{train}=25\%$ and $r_{train}=50\%$ in Fig. 4. We discuss the details of creating this version of the dataset in Supplementary. We can see how, indeed, the baseline model performs suboptimally, similarly as in Ego4D-AR in Sec. 4.4. Using random-replace with $p=60\%$ significantly improves the baseline accuracy at $r_{test}=100\%$ by 20 points when $r_{train}=25\%$ and 30 points when $r_{train}=50\%$. Furthermore, using modal-complete samples to train MMT ($p=60\%$) causes a significant performance boost compared to the model trained with modal-incomplete samples only ($p=0\%$).
Figure 5: Results on Epic-Sounds with $r_{train}^{A}=25\%,r_{train}^{V}=25\%$. We train our model with two MMTs: one for missing video and one for audio. We run the inference twice: (left) with missing video and (right) missing audio.
@@
Figure 6: Fusion Layer $L_{f}$ vs. accuracy in the models trained with no adaptation strategy (Baseline) and trained with MMT (Ours). For training with MMT, we use random-drop strategy with $p=60\%$ for Epic-Sounds and $p=25\%$ for Epic-Kitchens. This strategy makes MBT more robust to missing modalities across all $L_{f}$ and significantly reduces the negative effect of missing modalities.
4.6 Both modalities missing
In our setup, we train one MMT per dataset because we assume that only one modality could be missing. However, in more realistic scenarios, either modality could be missing for each sample in the dataset. Training a conventional multimodal baseline involves filtering out all modal-incomplete samples, causing the training data to shrink to a small intersection subset where all modalities are available, which could negatively affect the model’s performance.
Luckily, it is quite straightforward to extend our approach to two MMTs, one for each modality. To simulate this scenario, we use Epic-Sounds and remove audio from $r_{train}^{A}=25\%$  [^3] We use superscripts {A,V} to refer to the missing modality of the training samples and video from another $r_{train}^{V}=25\%$ of the training samples. For simplicity, we set $p=0$ for both MMTs when training our adapted model. To train a multimodal baseline, we filter out all modal-incomplete samples i.e., half of the training samples. Fig. 5, we show how our model’s performance compares to the baseline. We evaluate it for missing video (left) and audio (right). As can be seen, training with MMT dramatically improves the model performance.
@@ 4.6 Both modalities missing
As we mentioned in Sec. 3.2, the fusion layer does affect the performance of the bottleneck model, especially when test inputs are modal-incomplete. In Fig. 6, we examine whether training with MMT enhances robustness to a missing modality at various fusion layers and whether models trained with MMT (ours) show the same sensitivity to the fusion layer as those trained without it (baseline). As we observe, the introduction of MMT makes the model more robust to the missing modality across all fusion layers in both Epic-Sounds and Epic-Kitchens. With extremely severe $r_{test}=100\%$, the adapted models perform with $\sim 45\%$ accuracy in Epic-Kitchens, while the unimodal audio model achieves $40\%$ in this dataset. Furthermore, these adapted models do not exhibit similar sensitivity to the fusion layer as the baseline. For example, in Epic-Sounds, the baseline models trained with $L_{f}=11$ exhibit superior performance, which is not the case for the adapted models (at $r_{test}=100\%$, $37\%$ accuracy with $L_{f}=11$ but $\sim 40\%$ for other $L_{f}$). Overall, all adapted models perform consistently well, each providing decent performance in a modal-incomplete inference. Our approach effectively addresses a long-standing issue [35] of selecting the appropriate fusion layer when faced with missing modalities.
4.8 MMT vs. other missing modality representations
In Tables 2, we report the results using our random-replace and the baselines from Sec. 3.3 for Epic-Sounds, Epic-Kitchens, and Ego4D-AR. We also report the accuracy of the unimodal model in each dataset. Interestingly, passing zeroes as missing modality works better for Epic-Sounds and skipping missing modality tokens works better for Epic-Kitchens and Ego4D-AR.
@@
To compare to the baseline tailored to the missing modality problem, we implement the multimodal prompts method [30]. It was originally based on ViLT [27] for image-text classification, so we implemented our version for action recognition based on MBT. Note that this method relies on a strong pre-trained backbone to efficiently finetune it by optimizing a few network parameters. As we deal with audiovisual learning in egocentric videos, we rely on large-scale egocentric pre-training using MAEs. As seen in Tab. 2, our method outperforms the one proposed by [30].
@@
Table 2: Comparison of our method with baselines in Epic-Sounds, Epic-Kitchnes, and Ego4D-AR. We demonstrate the accuracy across different missing modality ratios $r_{test}$. We show in bold the best result and underline the runner-up. We mentioned the way of representing the missing modality in brackets: zeros for the baseline of passing zeros, skip for the baseline skipping the tokens of the missing input, as mentioned in Sec. 3.3, or MMT.
Dataset $r_{test}$ Unimodal Baseline Baseline Prompts [30] Ours
(zeros) (skip) (MMT)
Epic-Sounds 0% 41.4 55.2 55.2 36.7 56.3
25% 41.4 45.6 47.0 33.3 52.3
50% 41.4 37.1 39.9 30.0 48.6
75% 41.4 28.3 32.5 26.1 44.6
100% 41.4 19.5 25.0 22.6 40.7
Epic-Kitchens 0% 40.0 63.9 63.9 32.5 63.4
25% 40.0 55.5 53.2 31.1 59.0
50% 40.0 46.8 42.1 30.2 54.6
75% 40.0 37.9 30.8 29.2 50.0
100% 40.0 29.5 20.0 28.2 45.3
Ego4D-AR 27% 34.6 30.5 32.5 18.4 36.7
50% 34.6 25.9 30.8 16.4 35.2
75% 34.6 21.1 29.1 13.9 33.5
100% 34.6 16.0 27.2 11.5 32.0
@@
We find that across all datasets, our MMT trained with random-replace either reach (in Epic-Sounds and Ego4D-AR) or exceed (in Epic-Kitchens) the unimodal performance in extreme $r_{test}=100\%$. Interestingly, in Epic-Sounds, random-replace also regularizes the training and increases the $r_{test}=0\%$ performance by 1.1 points.
@@ 5 Conclusion
We explore the missing modality problem in multimodal egocentric datasets. We suggest a simple yet effective method by learning the optimal token representation of the missing modality (MMT). Placing learnable tokens to represent missing inputs provides an easy and intuitive way to train and test with modal-incomplete inputs. We propose strategy random-replace to learn MMT when training action recognition models and show how their performance brings us closer to robust and effective multimodal systems.